# Cost-Efficient Subjective Task Annotation and Modeling through Few-Shot Annotator Adaptation

### v1

@@ 5.1 Few Annotators are Enough
Ours $k=16$ 0.422(0.012) 0.44(0.025) 0.455(0.007) 0.795(0.009) 0.79(0.006) 0.785(0.005)
Ours $k=32$ 0.428(0.006) 0.447(0.019) 0.452(0.016) 0.795(0.01) 0.791(0.006) 0.785(0.004)
Ours $k=64$ 0.433(0.012) 0.451(0.015) 0.456(0.013) 0.797(0.009) 0.791(0.006) 0.785(0.004)
Ours $k=128$ 0.439(0.015) 0.457(0.012) 0.455(0.011) 0.798(0.008) 0.791(0.005) 0.786(0.004)
Table 1: Mean and standard deviation of $F_{1}$ scores on Brexit and MFSC dataset across three runs, with varying annotation budgets (%$\%B$). The budget allocation differs in the baseline and our framework; For example, $50\%\lvert D_{a_{i}}\rvert$ indicates selecting 50% of each annotator’s data, while $50\%\lvert\mathcal{A}\rvert$ denotes selecting $50\%$ of annotators and using all their data (i.e., 3 out of 6 annotators for Brexit, and 12 out of 24 annotators for MFSC).
Previous state-of-the-art methods for subjective tasks achieve performance gains by jointly learning annotations from multiple annotators, as opposed to independently learning each annotator’s perspectives Davani et al. (2022). In this analysis, we further explore whether adding more annotators to the multitask model results in a better model for all annotators. To answer this question, we look at the multitask performance $F^{mtl}_{1}$ as the number of annotators in the multitask model, $|\mathcal{A}_{mtl}|$, grows. There is no statistically significant correlation between the number of annotators and $F_{1}^{mtl}$ ($r=-0.28,p>0.05$). Figure 2 shows that average $F^{mtl}_{1}$ for all annotators does not increase as the number of annotators increases. Note that this pattern is consistent in both datasets and as Figure 2 shows, the $F^{mtl}_{1}$ scores are reliable with low standard deviations across models with different numbers of annotators. This observation motivates our first design choice; in the first stage of our framework, we can only rely on a small number of annotators to get a reliable multitask model.
Figure 2: Multitask models $F_{1}$ performance ($F_{1}^{mtl}$) as the number of annotators increase. Mean and standard deviation are reported across three repeated runs
5.2 Towards Better Performance with Less Annotation Budget
A successful framework for modeling subjective tasks should, above all, demonstrate the ability to accurately predict labels from all of the annotators. Therefore, we assess the overall performance ($F^{\text{overall}}_{1}$) of our proposed framework in comparison to the baseline of multitask learning. Recall the second component of our framework is to augment the multitask learning model of the first stage $\mathcal{F}_{mtl}$ to predict new annotators’ labels using only a few samples. To isolate the impact of the few-shot learning model from the specific choice of samples, we sample the $k$ shots randomly.
As shown in Table 1, our framework outperforms the multitask baseline on both datasets when it comes to predicting labels of all annotators. Specifically, on MFSC, our approach outperforms multitask learning using 100% of data with access to as little as 25% of the original annotation budget. On Brexit, our framework matches the performance of multitask learning on 100% budget with access to only 50% of the original budget. The findings demonstrate the success of our framework in achieving its dual objectives: enhancing performance across all annotators while reducing annotation budget requirements. Consequently, our framework facilitates increased diversity by incorporating additional annotator perspectives while adhering to a specific budget constraint. We also conduct an ablation study by omitting the first MTL stage and employing random few-shot sampling for each annotator (Appendix B.1).
$metric=F_{1}^{fs}\uparrow$ Brexit MFSC
$50\%$ $66\%$ $83\%$ $100\%$ $25\%$ $50\%$ $75\%$ $100\%$
$k=16$ $\mathcal{S}_{bal}$ 0.442(0.003) 0.457(0.016) 0.458(0.044) 0.785(0.006) 0.781(0.004) 0.778(0.01)
$\mathcal{S}_{dis}$ 0.399(0.013) 0.412(0.006) 0.457(0.036) 0.793(0.01) 0.795(0.004) 0.785(0.01)
$\mathcal{S}_{mv}$ 0.409(0.018) 0.416(0.026) 0.442(0.055) 0.799(0.003) 0.794(0.004) 0.786(0.015)
$\mathcal{S}_{rand}$ 0.4(0.017) 0.414(0.045) 0.44(0.037) 0.815(0.026) 0.814(0.033) 0.805(0.033)
$k=32$ $\mathcal{S}_{bal}$ 0.454(0.013) 0.469(0.004) 0.461(0.026) 0.787(0.006) 0.782(0.003) 0.779(0.012)
$\mathcal{S}_{dis}$ 0.402(0.008) 0.413(0.02) 0.455(0.03) 0.796(0.01) 0.797(0.004) 0.787(0.011)
$\mathcal{S}_{mv}$ 0.404(0.029) 0.426(0.029) 0.461(0.024) 0.801(0.003) 0.795(0.003) 0.787(0.017)
$\mathcal{S}_{rand}$ 0.412(0.006) 0.436(0.026) 0.422(0.027) 0.816(0.028) 0.815(0.035) 0.807(0.031)
$k=64$ $\mathcal{S}_{bal}$ 0.462(0.01) 0.47(0.018) 0.467(0.011) 0.788(0.005) 0.784(0.001) 0.782(0.011)
$\mathcal{S}_{dis}$ 0.429(0.018) 0.458(0.018) 0.518(0.017) 0.797(0.011) 0.799(0.005) 0.788(0.012)
$\mathcal{S}_{mv}$ 0.409(0.017) 0.41(0.05) 0.467(0.011) 0.802(0.004) 0.797(0.002) 0.79(0.018)
$\mathcal{S}_{rand}$ 0.423(0.02) 0.447(0.016) 0.447(0.019) 0.818(0.027) 0.815(0.033) 0.809(0.033)
$k=128$ $\mathcal{S}_{bal}$ 0.498(0.009) 0.517(0.025) 0.524(0.021) 0.79(0.005) 0.785(0.001) 0.785(0.01)
$\mathcal{S}_{dis}$ 0.456(0.011) 0.476(0.026) 0.511(0.023) 0.798(0.009) 0.799(0.003) 0.786(0.012)
$\mathcal{S}_{mv}$ 0.425(0.028) 0.428(0.035) 0.46(0.025) 0.805(0.004) 0.798(0.003) 0.79(0.019)
$\mathcal{S}_{rand}$ 0.433(0.023) 0.466(0.006) 0.443(0.027) 0.819(0.026) 0.816(0.032) 0.81(0.031)
Table 2: Few-shot $F_{1}$ results ($F_{1}^{fs}$) on BREXIT and MFSC datasets across varying percentages of the full budget (%$\%B$). Mean and standard deviation calculated over three repeated runs.
5.3 Few-shot Sampling Strategies
Recall that in the second stage of our framework, we only select a small subset of the input texts $\mathcal{S}(X)\subset X$ to be annotated by the new annotators (Figure 1). Intuitively, we are relying on the first stage of our framework (i.e., multitask learning) to learn the commonalities between annotators. Hence, the goal of the selection function $\mathcal{S}(X)$ is to choose input texts that best capture individual annotators’ differences. To study the impact of data selection strategies, we compare the fewshot performance ($F^{fs}_{1}$) of each strategy controlling for the number of shots $k$ (see §3 for definitions of $\mathcal{S(.)}$).
On Brexit dataset, the $\mathcal{S}_{bal}$ strategy, which selects a balanced sample for each annotator outperforms other approaches consistently with 50% budget. With 66% budget, $\mathcal{S}_{bal}$ strategy performs the best for all values of $k$. Notably, with this budget, the performance gap is closed by the $\mathcal{S}_{dis}$ for higher number of shots. Although the differences in performance are less pronounced for higher budget scenarios (i.e., the 83%$B$), $\mathcal{S}_{bal}$ and $\mathcal{S}_{dis}$ still achieve better results especially when more shots are available. We acknowledge that the $\mathcal{S}_{bal}$ strategy relies on access to a balanced sample for each annotator which is often not readily available. However, this study showcases the potential of this approach and future work can explore methods that attempt to select instances that approximate a balanced sample per annotator.
On the MFSC dataset, sampling strategies persistently outperform the multitask learning baseline in all scenarios. While different strategies seem to perform comparably, a crucial differentiating factor is the impact of various strategies on the stability of the results. Specifically, compared to $\mathcal{S}_{rand}$, all sampling functions result in more stable performances (i.e., lower standard deviation) across all budgets and values of $k$.
Overall, our results on two datasets show that the fewshot stage of our framework, results in models that consistently outperform the multitask learning baseline. Our experiments clearly demonstrate the importance of sample selection strategies both in terms of performance improvement and stability. Furthermore, our results on both datasets motivate the exploration of sampling strategies that can approximate a balanced sample per annotator.
5.4 Reduced Performance Disparities across Annotators
Ensuring a comprehensive representation of annotators’ viewpoints is crucial in modeling subjective tasks. To achieve this goal, a critical criterion is to create models that not only improve the aggregated performance but also demonstrate fair and equitable performance across all annotators. For example, if the $F_{1}$ scores of one model for two annotators are 0.6 and 0.8, respectively, while the second model scores 0.7 for both annotators, the latter is considered a better model. Although the average performance is the same for both models, the first model has a disparate negative impact on the first annotator.
@@ 5.4 Reduced Performance Disparities across Annotators
Figure 3: Comparison of Annotator level F1 scores ($F_{1}^{a_{i}}$) on the Brexit dataset between MTL model and our framework, leveraging the $\mathcal{S}_{bal}$ sampling method for all budgets and shots
This is important because performance disparities among social groups (in our case annotators) can lead to biased models, limiting the system’s ability to accurately reflect diverse perspectives and potentially perpetuating inequalities in the outputs of subjective tasks (Buolamwini and Gebru, 2018). Merely relying on aggregated performance measures, like the average across all annotators, falls short of providing a complete understanding of how well the model is capturing annotators’ varying perspectives. For example, it is not clear whether the average performance is improving because an approach improves on capturing only a subset of annotators or for everyone. Hence, we look into the standard deviation of performance across all annotators $d=\sqrt{\frac{1}{N-1}\sum_{i=1}^{N}(F^{a_{i}}_{1}-\overline{F^{overall}_{1}})^{2}}$. Lower standard deviations are indicative of more equitable models. As shown in Table 3, regardless of the sampling strategy, our approach results in lower performance variances compared to the MTL baseline on lower budgets (50% and 66%). Comparing the different sampling strategies on Brexit dataset, we observe that $S_{bal}$ achieves the most equitable models across all values of $k$. On MFSC, using a balanced sampling strategy consistently reduces the disparities among annotators compared to the MTL baseline for all budgets. Figure 3 visualizes the performance of our framework in comparison to the multitask learning baseline for each annotator. Notably, our framework improves the performance for the annotators in the non-Western control group (i.e., the first three annotators) while maintaining the performance of the rest of the annotators.
Overall, these results suggest that our proposed framework not only improves the overall performance of all annotators but also yields models that are more fair and equitable. Interestingly, we observe similar patterns in fairness and overall performance improvements. In other words, the configurations with the best overall performance (see section 5.3) are also the ones that have the least performance disparities among all annotators.
5.5 Annotator level analysis
@@
Here we delve into the relationship between annotator-level variables. Recall that our framework is trained on $\mathcal{A}_{mtl}$ in the initial stage, followed by fine-tuning for each $a\in\mathcal{A}_{fs}$. Hence, a practical question arises: does the choice of the set $\mathcal{A}_{mtl}$ matter? In other words, would the similarity or divergence in perspectives among annotators in this set impact the performance on $\mathcal{A}_{fs}$? Investigating this is crucial, as identifying such an effect would necessitate a thoughtful selection of $\mathcal{A}_{mtl}$. To examine this, we conduct the following analysis:
  $metric=d\downarrow$ Brexit MFSC
$50\%$ $66\%$ $83\%$ $25\%$ $50\%$ $75\%$
  MTL .168 .139 .131 .128 .136 .127
$k=16$ $\mathcal{S}_{bal}$119 .127 .132 .119 .124 .124
$\mathcal{S}_{dis}$141 .134 .127 .129 .130 .126
$\mathcal{S}_{mv}$138 .137 .131 .131 .130 .127
$\mathcal{S}_{rand}$137 .139 .135 .137 .135 .130
$k=32$ $\mathcal{S}_{bal}$113 .125 .133 .120 .125 .124
$\mathcal{S}_{dis}$145 .138 .131 .131 .130 .126
$\mathcal{S}_{mv}$140 .137 .129 .130 .130 .127
$\mathcal{S}_{rand}$139 .135 .138 .136 .134 .129
$k=64$ $\mathcal{S}_{bal}$109 .119 .127 .119 .124 .123
$\mathcal{S}_{dis}$137 .126 .123 .131 .130 .126
$\mathcal{S}_{mv}$145 .146 .126 .130 .129 .127
$\mathcal{S}_{rand}$131 .128 .132 .136 .135 .129
$k=128$ $\mathcal{S}_{bal}$093 .108 .117 .117 .122 .121
$\mathcal{S}_{dis}$111 .120 .124 .130 .129 .127
$\mathcal{S}_{mv}$137 .142 .132 .126 .128 .127
$\mathcal{S}_{rand}$131 .127 .136 .134 .133 .128
[…]
@@
Disagreement within $\mathcal{A}_{mtl}$ and performance on $\mathcal{A}_{fs}$: The aim of this analysis is to investigate whether there is a relationship between the disagreement within annotators in $\mathcal{A}_{mtl}$ and the performance of the newly adopted annotators in $A_{fs}$. To test this relationship, we employ a mixed-effects model to predict the performance of $a\in\mathcal{A}_{fs}$ by the agreement within $\mathcal{A}_{mtl}$ denoted as $d^{1}$ (Fleiss, 1971). The model controls for $k$, budget $B$, and agreement between $\mathcal{A}_{fs}$ and $\mathcal{A}_{mtl}$, denoted using $d^{2}$, incorporating random effects for $\mathcal{A}_{mtl}$ and $\mathcal{A}_{fs}$. The formula for this model is as follows:
@@
The findings regarding Brexit indicate that there is no statistically significant effect of agreement within $\mathcal{A}_{mtl}$ ($d^{1}$) on the performance. For the MFSC dataset, a significant effect was observed only for results obtained from $\mathcal{S}_{bal}$ ($\beta_{1}=-0.052,SE=0.012,p<0.001$). This implies that a unit decrease in $d^{1}$, corresponding to moving from full agreement to full disagreement, is associated with a 0.052 increase in the $F_{1}$ score. This finding suggests that selecting a diverse $\mathcal{A}_{mtl}$ with high disagreement can potentially be advantageous.
@@ 6 Conclusion
We introduced a framework for annotation collection and annotator modeling in subjective tasks. Our framework aims to minimize the annotation budget required to model a fixed number of annotators while maximizing the predictive performance for each annotator. Our approach involves collecting annotations from an initial set of annotators and building a multitask model that captures general task patterns while signaling differences among individual annotators. Subsequently, we utilize the annotations from the first stage to select a small set of samples from new annotators that best highlight their deviations from the general patterns. Finally, we leverage this collected data to augment the initial model, enabling it to learn the new annotator’s perspective from a limited number of examples. We explored four distinct methods for few-shot sample selection and found that the most effective approach involves balanced sample selection. We introduced a new subjective task dataset Moral Foundations Subjective Corpus (MFSC ), of 2000 Reddit posts annotated by 24 annotators for moral sentiment which enabled us to test our framework in scale. Our experiments on MFSC and a hate speech dataset revealed that our framework consistently surpasses previous SOTA with access to as little as 25% of the original annotation budget. In addition, we showed that our framework yields more equitable models that reduce performance disparities among annotators. Our cost-effective framework for subjective tasks allows increasing the number of annotators, enhancing the diversity of perspectives captured, and facilitates the integration of new annotators into pre-existing datasets.
@@ Appendix A Dataset details
Table 8 displays detailed information about each dataset, including examples and annotations of all annotators.
@@ A.1 Demographics of MFSC Annotators
Appendix B Experiments and Results
B.1 Additional Baseline and Ablation Study
First, we conduct an ablation study where we omit the first stage MTL. Essentially, this model is equivalent to few-shot adaptation for each annotator initializing from a pre-trained model. The resulting $F_{1}$ scores are presented in Table 4. Comparing these scores with our framework using random few-shot sampling for $k=16$ from Table 2, we observe that our framework has a 21% gain for Brexit and a 15.5% gain for MFSC. This study demonstrates that the first stage MTL in our framework is crucial for its success.
In the second ablation study, we omit the second stage few-shot sample selection from our framework. In other words, in the second stage, we used the entire annotated samples for each annotator instead of selecting only a few samples. Note that this is equivalent to using 100% of the budget and serves as an upper bound to the performance achieved with an ideal sampling function. Additionally, we present Ensemble, a new baseline in which we train a separate model for each annotator using 100% of data from that annotator. Figure 5 presents a comparison of 3 different strategies, using 100% of the budget (MTL, Ensemble, and ours). On the Brexit Dataset (top) our framework has as much as 7.4% performance gain compared to Ensemble baseline (when using $\frac{4}{6}$ annotators in MTL), and for MFSC dataset our framework has as large as 5% gain compared to Ensemble baseline (when using $\frac{12}{24}$ of annotators in MTL). These results show that even considering the 100% budget our framework outperforms both baselines demonstrating the benefit of our two-stage design.
Figure 5: Comparision of our framework with baselines for the two datasets
Dataset $k=16$ $k=32$ $k=64$ $k=128$
Brexit 0.23 0.22 0.20 0.28
MFSC 0.66 0.73 0.71 0.77
Table 4: $F_{1}$ scores for few-shot baseline with random sampling $\mathcal{S}_{rand}$ when the first stage MTL is omitted.
B.2 Individual Moral Concern Results

### v2

@@ 5.2 Reduced Performance Disparities across Annotators
  Roberta-Base MTL .168 .139 .131 .130 .128 .136 .127 .130
$S_{bal}$093 .108 .117 .117 .122 .121
$S_{dis}$111 .120 .124 .130 .129 .127
$S_{mv}$137 .142 .132 .126 .128 .127
$S_{rand}$131 .127 .136 .134 .133 .128
Roberta-Large MTL .170 .136 .117 .155 .152 .143 .146 .149
$S_{bal}$102 .127 .148 .117 .121 .127
$S_{dis}$112 .134 .140 .132 .128 .133
$S_{mv}$134 .141 .149 .130 .128 .131
$S_{rand}$120 .131 .146 .127 .129 .130
Llama-3 MTL .172 .122 .112 .130 .189 .171 .167 .158
$S_{bal}$117 .113 .120 .137 .147 .158
$S_{dis}$114 .113 .120 .198 .187 .176
$S_{mv}$127 .103 .124 .170 .174 .170
$S_{rand}$150 .128 .119 .176 .173 .166
Table 3: Performance disparities across annotators ($d\downarrow$). The best values are shown in bold. $\mathcal{S}_{bal}$, $\mathcal{S}_{dis}$, $\mathcal{S}_{mv}$, and $\mathcal{S}_{rand}$ refer to the sampling functions used in the second stage of our framework (§3). We generally observe lower performance disparities with our framework compared to the MTL baseline.
@@
Figure 4: Comparison of Annotator level $F_{1}$ scores ($F_{1}^{a_{i}}$) on the Brexit dataset between MTL model and our framework, leveraging the $\mathcal{S}_{bal}$ sampling method for all budgets and shots on RoBERTa-base model
5.3 Annotator-level Analysis
@@
Here, we delve into the relationship between annotator-level variables. Recall that our framework is trained on $\mathcal{A}_{mtl}$ in the initial stage, followed by fine-tuning for each $a\in\mathcal{A}_{fs}$. Hence, a practical question arises: does the choice of the set $\mathcal{A}_{mtl}$ matter? In other words, would the similarity or divergence in perspectives among annotators in this set impact the performance on $\mathcal{A}_{fs}$? Investigating this is crucial, as identifying such an effect would necessitate a thoughtful selection of $\mathcal{A}_{mtl}$. To examine this, we conduct the following analysis:
@@
Disagreement within $\mathcal{A}_{mtl}$ and performance on $\mathcal{A}_{fs}$: The aim of this analysis is to investigate whether there is a relationship between the disagreement within annotators in $\mathcal{A}_{mtl}$ and the performance of the newly adopted annotators in $A_{fs}$.
To test this relationship, we employ a mixed-effects model to predict the performance of $a\in\mathcal{A}_{fs}$ by the agreement within $\mathcal{A}_{mtl}$ denoted as $d^{1}$ (Fleiss, 1971). The model controls for $k$, budget $B$, and agreement between $\mathcal{A}_{fs}$ and $\mathcal{A}_{mtl}$, denoted using $d^{2}$, incorporating random effects for $\mathcal{A}_{mtl}$ and $\mathcal{A}_{fs}$. The formula for this model is as follows:
@@
The findings regarding Brexit indicate no statistically significant effect of agreement within $\mathcal{A}_{mtl}$ ($d^{1}$) on the performance. For the MFSC dataset, a significant effect was observed only for results obtained from $\mathcal{S}_{bal}$ ($\beta_{1}=-0.052,SE=0.012,p<0.001$). This implies that a unit decrease in $d^{1}$, corresponding to moving from full agreement to full disagreement, is associated with a 0.052 increase in the $F_{1}$ score. This finding suggests that selecting a diverse $\mathcal{A}_{mtl}$ with high disagreement can potentially be advantageous.
@@ 6 Conclusion
We introduced a framework for annotation collection and annotator modeling in subjective tasks. Our framework aims to minimize the annotation budget required to model a fixed number of annotators while maximizing the predictive performance for each annotator. Our approach involves collecting annotations from an initial set of annotators and building a multitask model that captures general task patterns while signaling differences among individual annotators. Subsequently, we utilize the annotations from the first stage to select a small set of samples from new annotators that best highlight their deviations from the general patterns. Finally, we use these samples to augment the initial model in a few-shot setting to learn the new annotator’s perspective. We evaluated our framework using three base models, and explored four distinct methods for few-shot sample selection and found that the most effective approach involves balanced and random sample selections. We introduced a new subjective task dataset Moral Foundations Subjective Corpus (MFSC), of 2000 Reddit posts annotated by 24 annotators for moral sentiment which enabled us to test our framework in scale. Our experiments on MFSC and a hate speech dataset revealed that our framework consistently surpasses previous SOTA with access to as little as 25% of the original annotation budget. In addition, we showed that our framework yields more equitable models that reduce performance disparities among annotators. Our cost-efficient framework for subjective tasks allows enhancing the diversity of the captured perspectives, and facilitates the integration of new annotators into pre-existing datasets and models.
@@ Appendix A Dataset Details
Each sample in the MFSC dataset is annotated with a binary label indicating moral sentiment: 1 if the sentence pertains to morality and 0 if it does not. Additionally, we have collected more fine-grained labels of morality (i.e., Purity, Harm, Loyalty, Authority, Proportionality, or Equality) following the procedure outlined by Trager et al. (2022). Specifically, if a sample is labeled as moral, the annotator can select any of the applicable moral categories for that text. The distribution of these labels across annotators is demonstrated in Figure 6. The dataset also includes additional metadata information, such as confidence levels for each instance using a 3-level measure (confident, somewhat confident, and not confident). Furthermore, we collected annotator responses for the “Big Five Inventory-2-Short” questionnaire (Soto and John, 2017). The MFSC dataset provides an opportunity to explore the subjective nature of morality. The substantial number of annotators, along with their questionnaire responses, enables future researchers to investigate the modeling of subjective tasks on a larger scale. See Table 1 and Table 4 for sample annotations for MFSC and Brexit datasets.
Brexit examples a1 a2 a3 a4 a5 a6
THE MAJORITY WILL NEVER allow the Mentally Ill Globalists to turn the world into a SJW and Radical Islam SAFE SPACE #brexit #Trump2016 0 0 1 1 0 1
A muslim Mayor of London? What!? This PC Sickness has become a pandemic. England turning into Little Asia. 0 0 0 1 1 1
Not all foreign people who wants to go to the uk have bad intentions. Improve your law. The #Brexit isn’t gonna help your economy. 0 0 0 0 0 0
Table 4: Examples from Brexit dataset with binary Hate labels from all 6 annotators.
@@ A.1 Demographics of MFSC Annotators
Figure 6: Distribution of the labels across annotators in MFSC dataset
Appendix B Additional Baseline and Ablation Study
In the following sections, we conduct the experiments using the Roberta-base model due to its superior performance among the base models in our experiments, as well as its resource efficiency.
First, we conduct an ablation study by omitting the first stage of MTL, effectively reducing the model to few-shot adaptation for each annotator from a pre-trained model. The resulting $F_{1}$ scores are shown in Figure 8. When comparing these scores to our complete framework in Table 8, we observe that our framework consistently outperforms the second stage alone in all few-shot scenarios. For instance, in random few-shot sampling for $k=16$, our model achieves a 23% gain in Brexit and a 12% gain in MFSC compared to this ablation model. This highlights the critical role of the first stage of MTL in the success of our framework.
In the second ablation study, we omit the second stage, few-shot sample selection, from our framework. In other words, in the second stage, we use all annotated samples for each annotator instead of selecting only a few samples. Note that this is equivalent to using 100% of the budget and serves as an upper bound to the performance achieved with an ideal sampling function.
Additionally, we present a new baseline where a separate model is trained for each annotator using 100% of their respective data. Following the naming convention used by Davani et al. (2022), we refer to this baseline as “Ensemble” to ensure consistency with previous work in this field.The Ensemble baseline involves fine-tuning the model directly for each annotator, calculating individual annotator $F_{1}$ scores, and reporting the average $F_{1}$ score across annotators. Hyperparameters and epoch numbers for training are consistent with those mentioned for the MTL model in Section 4.3. Figure 8 presents a comparison of 3 different strategies, using 100% of the budget (MTL, Ensemble, and ours). On the Brexit Dataset (top) our framework has as much as 7.4% performance gain compared to the Ensemble baseline (when using $\frac{4}{6}$ annotators in MTL), and for the MFSC dataset our framework has as large as 5% gain compared to Ensemble baseline (when using $\frac{12}{24}$ of annotators in MTL). These results show that even considering the 100% budget, our framework outperforms both baselines, demonstrating the benefit of our two-stage design. Interestingly, the Ensemble model outperforms MTL for these datasets, contrary to previous research findings comparing these two methods.
(a) Brexit
(b) MFSC
Figure 8: Baseline results in Blue compared to our framework results in green
Appendix C Additional Tasks
C.1 MFSC (Care label)